OpenStudy (kkutie7):

Anti derivative using u substitution for \[\frac{6x}{(1+x^{2})^{2}\]

OpenStudy (kkutie7):

\[\frac{6x}{(1+x^{2})^{2}}\]

jimthompson5910 (jim_thompson5910):

let u = 1+x^2 deriving both sides with respect to x gives du/dx = 2x du = 2x*dx du/2 = x*dx x*dx = du/2 6*x*dx = 6*du/2 6x*dx = 3du so we go from \[\int \frac{6x*dx}{(1+x^{2})^{2}}\] to \[\int \frac{3du}{u^{2}}\] \[3\int \frac{1}{u^{2}}*du\]

jimthompson5910 (jim_thompson5910):

do you see how to finish up?

OpenStudy (kkutie7):

umm give me a min.

jimthompson5910 (jim_thompson5910):

ok

jimthompson5910 (jim_thompson5910):

if you are stuck, then keep in mind that \[\Large \frac{1}{u^{2}} = u^{-2}\]

OpenStudy (kkutie7):

so \[-\frac{3}{1+x^{2}}\]

jimthompson5910 (jim_thompson5910):

don't forget the +C

OpenStudy (kkutie7):

oh I don't need the constant for the problem I'm working on but that you =)

OpenStudy (kkutie7):

*thank

jimthompson5910 (jim_thompson5910):

otherwise, \[\int \frac{6x*dx}{(1+x^{2})^{2}} = -\frac{3}{1+x^2}+C\] is correct. Nice work

jimthompson5910 (jim_thompson5910):

oh you're doing a definite integral (area under the curve), I see

OpenStudy (kkutie7):

yupp

jimthompson5910 (jim_thompson5910):

you're welcome